OpenStudy (anonymous):

When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week? ____mi

OpenStudy (anonymous):

\[40+x=4x + 10\] Notice the key language here: = the same as + added to + adding * times the number. Now isolate the x on the left side of the equation and move everything else over to the right side.

OpenStudy (anonymous):

\[40-10=4x-x\] \[30=3x\] \[\frac{ 30 }{ 3 }=x\] \[x=10\]

OpenStudy (anonymous):

double check your work by putting 10 back in to your original equation: \[40+10 = 4(10) + 10\] \[50=40+10\] \[50=50\] :)

OpenStudy (anonymous):

ohh I get it now :) thank you

OpenStudy (anonymous):

no problem, word problems are the bane of my existence so happy i could help for once :D

OpenStudy (anonymous):

yeah im not good with word problems haha